OpenStudy (anonymous):

Wats the derivative of pi^2?

OpenStudy (adamaero):

the derivative of any constant is zero

OpenStudy (anonymous):

but it's 2 pi

OpenStudy (anonymous):

\[ \frac{d}{dx}x^2\Bigg|^{x=\pi} = 2\pi \]However: \[ \frac{d}{dx}\pi^2 = 0 \]

OpenStudy (anonymous):

or maybe its like two pie , it hast to be stated out